OpenStudy (fairegaming):

**How many medals can I get on one post?** Also Answering Questions if I know them!

OpenStudy (zzr0ck3r):

any natural number n

OpenStudy (zzr0ck3r):

or 0

OpenStudy (anonymous):

ANY NUMBER

OpenStudy (zzr0ck3r):

if you think 0 is natural then -1

OpenStudy (anonymous):

CAN SOMEBODY HELP ME FAST

OpenStudy (zzr0ck3r):

tp be fair @kadijahbaby34334 you cant get \(\sqrt{2}\) medals

OpenStudy (anonymous):

WHAT

OpenStudy (fairegaming):

I sound like a d*** but can I help you guys 4 a medal? :P

OpenStudy (anonymous):

HELP

OpenStudy (fairegaming):

Question Please

OpenStudy (anonymous):

Pasha works as a food scientist. He designed a recipe for crackers that are 35% wheat flour by weight. He made a test batch of the crackers from 1 3/4 total pounds of ingredients. What weight of wheat flour did Pasha use in the test batch? A. 1 11/80 B. 1 2/5 C. 49/80 HURRY PLEASE IT IS TIMED

OpenStudy (anonymous):

NEVER MIND I GOT A 0%

OpenStudy (fairegaming):

:/

OpenStudy (aaronandyson):

1 3/4 = 7/4 7/4 * 40% = 7/4 * 4/10 = 28/40 = 14 / 20 = 7/10 pounds

iYuko (iyuko):

\(\infty\)

OpenStudy (anonymous):

thx every one